7eeST 3 LMahae FaD DecACTls CHeK (PLACE ECACTION A STRJUTUce 3Kt VIK) MK)
Answer #1

step 1: find out reactions by using equlibrium equations

step 2: Now cut the each segment and apply shear and moment equationsOn force AA 12! Kt 7SLft SED V,40 VA /A/e Ay AVe -ve 4vel SFy= Ay+ Cy= S6+40 Ay+Cy= 46RN Qb3 -40x24-75 0. -36x6 +yx 18 A8 Cy=) Cut Sec Betweon cEa Cos6) tus 36t AMCD 65+ GA.5-SC VeD 17 69-SN 6 SK (2( Vevo 76-36 10 McD 6.5 Ct)+ 69.52 36 C12+2) Me -315